OpenStudy (anonymous):

what quadrant is 5pi/2 in?

OpenStudy (anonymous):

0 to pi/2 is Q1, pi/2 to pi is Q2, pi to 3pi/2 is Q3 and 3pi/2 to 0 is Q4. Do the math

OpenStudy (anonymous):

5pi/2 = 2pi +pi/2 2pi makes one complete angle ...

OpenStudy (wolf1728):

(5 PI) / 4 = 225 (5 PI) / 2 = 450 450° is equivalent to 90° and would be between quadrants 1 and 2